Let f: [-5,5]-->R be defined by f(x) =5[x] +x^2 , where [x] denotes the greatest integer function . Show that this function is integrable .Is this function also differentiable ? Justify your answer ?
1
Expert's answer
2019-10-01T11:06:22-0400

"\int\limits_{-5}^5f(x)dx=\int\limits_{-5}^5x^2dx+5\int\limits_{-5}^5[x]dx". The integral "\int\limits_{-5}^5x^2dx" exists. Show that integral "\int\limits_{-5}^5[x]dx" also exists.

"\int\limits_{-5}^5[x]dx=\sum\limits_{k=-5}^4\int\limits_k^{k+1}[x]dx=\sum\limits_{k=-5}^4\int\limits_k^{k+1}kdx" , so integral "\int\limits_{-5}^5[x]dx" exists.

"f" is not differentiable in "[-4,5]\cap\mathbb Z", because left derivative of "f" in this set do not exist .

Indeed, let "k\in[-4,5]\cap\mathbb Z" , so left derivative in point "k" is"\lim\limits_{x\to 0+}\frac{f(k)-f(k-x)}{x}=\lim\limits_{x\to 0+}\frac{5[k]+k^2-5[k-x]-(k-x)^2}{x}="

"=\lim\limits_{x\to 0+}\frac{5k+k^2-5[k-x]-k^2+2kx-x^2}{x}=\lim\limits_{x\to 0+}\frac{5k-5[k-x]+2kx-x^2}{x}="

"=2k+5\lim\limits_{x\to 0+}\frac{k-[k-x]}{x}"

If "x\in (0,1)" , then "[k-x]=k-1" , so "\lim\limits_{x\to 0+}\frac{k-[k-x]}{x}=\lim\limits_{x\to 0+}\frac{k-(k-1)}{x}=+\infty"

So we obtain that left derivatives of "f" in "[-4,5]\cap\mathbb Z" equal "+\infty" , that is, do not exist.

Answer: "f" is integrable, but not differentiable.